OpenStudy (anonymous):

Helpful steps,What is the empirical formula of a compound composed of 38.5 g potassium (K) and 7.88 g oxygen (O)?

OpenStudy (matt101):

The empirical formula tells you the ratio of the different atoms in the compound. This means we need to know NUMBERS. The question has given us MASSES, so we need to do a conversion to a number, in this case moles. Just divide the mass by the molar mass: moles of K = 38.5 / 39 = 1 moles of O = 7.88 / 16 = 0.5 I've rounded a bit to make things easier. So our formula now is: \[KO_{0.5}\] But we can't have a fraction in the empirical formula, so let's just multiple everything by 2 to keep the same ratio by get rid of the 0.5: \[K_{2} O\] And that's your answer! If you have any questions let me know!

OpenStudy (anonymous):

Thanks for your help!